OpenStudy (anonymous):

What is the y-intercept of the line that has a slope of -1/2 and passes through the point (2, 3)?

OpenStudy (godlovesme):

r u familiar with the point slope formula? :) @ms1231988

OpenStudy (anonymous):

the y-intercept is the point where the line crosses the y-axis

OpenStudy (anonymous):

y=1/2x+ ______ I understand y=mx+b i just need help with the last numbers..

OpenStudy (anonymous):

So find the a line that crosses through (2, 3) and add a slope of -½ and then look at where it crosses the y-axis

OpenStudy (anonymous):

no the point slope is like this (x-x1) = m(x-x1)

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

now im even more confused....

OpenStudy (anonymous):

so fill in the known numbers: y-(-3) = -1/2(x-(-2)

OpenStudy (anonymous):

2-(-3) = -1/2(3-(-2) ...?

OpenStudy (anonymous):

then you can just make it in slope intercept form!: y+3=-1/2x-1 y = -1/2x -4

OpenStudy (anonymous):

so the y intercept is -4

OpenStudy (anonymous):

sorry this stuff is pretty complicated and i'm probably not doing a good job at explaining it :P

OpenStudy (anonymous):

I just learned this on VLACS like yesterday, I understand it pretty well

OpenStudy (anonymous):

thats the whole question... so i out -4 in the box?

OpenStudy (anonymous):

yes but you can try it out yourself by graphing the equation y=-1/2x -4 and see is it has the point 2,3

OpenStudy (anonymous):

if you have any questions about how I came to this answer look at this: https://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/point-slope-form/e/converting_between_point_slope_and_slope_intercept